We use the t distribution to construct a confidence interval for the population mean when the underlying population standard deviation is not known. Under the assumption that the population is normally distributed, find tα/2,df for the following scenarios. (You may find it useful to reference the t table. Round your answers to 3 decimal places.)

tα/2,df
a. A 90% confidence level and a sample of 25 observations.
b. A 95% confidence level and a sample of 25 observations.
Answer #1

solution

given that

(A)n = 25

Degrees of freedom = df = n - 1 = 25- 1 = 24

At 90% confidence level the t is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.1

\alpha / 2 = 0.1 / 2 = 0.05

t\alpha /2,df = t0.05,24 = 1.711 ( using student t table)

(B)

n = 25

Degrees of freedom = df = n - 1 = 25- 1 = 24

At 95% confidence level the t is ,

\alpha = 1 - 95% = 1 - 0.95 = 0.05

  \alpha/ 2= 0.05 / 2 = 0.025

t\alpha /2,df = t0.025,24 = 2.064 ( using student t table)